OpenStudy (anonymous):

Logarithms: 2^(x+1)=3^x

OpenStudy (tkhunny):

You should apply a logarithm. Since you have base 2 and 3, just pick a familiar base. \(log(2^{x+1}) = log(3^x)\) This leads to \((x+1)log(2) = xlog(3)\)

OpenStudy (anonymous):

umm so what is x = ?

OpenStudy (anonymous):

Solve for it.

OpenStudy (tkhunny):

This requires a little algebra. Don't let log(2) or log(3) scare you. Those are just numbers, like 7 or 4 or \(\sqrt{2}\) or \(\pi\).

OpenStudy (anonymous):

i got log2/ (log3-log2) not sure if its correct?

OpenStudy (tkhunny):

Well, there shoudl be an "x =" in there, somewhere. Good work.

OpenStudy (anonymous):

thanks